OpenStudy (anonymous):

@phi Could you please help me with this? Which logarithmic graph can be used to approximate the value of y in the equation 6y = 12

OpenStudy (anonymous):



OpenStudy (phi):

log(6^y) = log(12) or y log(6) = log(12) y = log(12)/ log(6)

OpenStudy (anonymous):

could you put it in a graphing way

OpenStudy (anonymous):

@phi are you there?

OpenStudy (anonymous):

What about this one... For the graphed function f(x) = (2)x + 2 + 1, calculate the average rate of change from x = -2 to x = 0.